OpenStudy (anonymous):

would e^ln(2x+3)= just 2x+3?

Parth (parthkohli):

yes.

OpenStudy (anonymous):

cool thanks :)

Parth (parthkohli):

is that all you wanted to ask?

OpenStudy (anonymous):

well im confused on log(10^3x^7)=?

Parth (parthkohli):

\[\log\left( 10^{3x^7}\right)\]this?

OpenStudy (anonymous):

\[\log({10^{3}} x ^{7})\] woops sorry

Parth (parthkohli):

oh.\[\log(10^3 x^7) = \log(10^3) + \log(x^7)\]

OpenStudy (anonymous):

does the problem solving continue from there?

Parth (parthkohli):

Yes.

OpenStudy (anonymous):

would the answer be \[3x ^{7}\]

OpenStudy (isaiah.feynman):

Yes

OpenStudy (isaiah.feynman):

You seem to know your logs well.

OpenStudy (anonymous):

lol thanks

OpenStudy (isaiah.feynman):

No. No sorry. I looked too quickly. It can't be 3x^7

OpenStudy (isaiah.feynman):

\[\log(10^{3})+\log (x^{7}) = 3+\log(x^{7})\]

OpenStudy (anonymous):

whys so?

OpenStudy (anonymous):

well would 150=50e^7x be x=ln(3)/7

OpenStudy (isaiah.feynman):

Yes.

OpenStudy (anonymous):

are you sure? lol

OpenStudy (isaiah.feynman):